At highway speeds, a particular automobile is capable of an acceleration of about 1.6 m/s2. At this rate, how long does it take to accelerate from 83 km/h to 108 km/h?
Answer #1

83km/hr = 83 * 5/18 m/s = 23.05 m/s
108 km/hr = 108*5/18 m/s = 30.0 m/s
v = u +at
v- u/a = t = 30.0-23.05/1.6 = 4.34s